Suppose labour and capital are perfectly substitutable at a 2:1 rate, in such a way that 2 units of labor or 1 unit of capital can produce 2 units of output. What would be the equation for the production function? What is the average and marginal product of labor in this case?
Answer #1

From the given information , we get the equation:

Total product ,Q = L +2K

Average product , Q/L = (L+2K)/ L = 1+2K/L

Marginal product of labor , dQ/dL = 1